Compute the extension in length of a wire 530 cm long and a diameter of 0.15 cm. When the body of mass (80*15) gm is applied vertically. The young modulus of given wire is 5x109 Dyne/cm2. 


1
Expert's answer
2021-06-22T09:47:46-0400

By the definition of the Young's modulus, we have:


"E=\dfrac{FL_0}{A\Delta L},""\Delta L=\dfrac{FL_0}{AE}=\dfrac{mgL_0}{\pi r^2E},""\Delta L=\dfrac{1.2\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot5.3\ m}{\pi(7.5\cdot10^{-4}\ m)^2\cdot5\cdot10^8\ \dfrac{N}{m^2}}=0.07\ m."